Estimate the sales for 2022, 2023, 2024 and fit a linear regression equation and draw a trend line.



Years 2014 2015 2016 2017 2018 2019 2020 2021


Sales 22734 24731 31489 44690 51320 55330 65890 75800



second question


The demand for petrol rises from 500 to 600 Barrels when the price of a particular scooter is reduced from ETB 25,000 to ETB 22,000. Find out the cross elasticity of demand for the two. What is the nature of their relationship?



1
Expert's answer
2022-03-30T14:13:03-0400

The trendline is as shown




Find out the cross elasticity of demand for the two. What is the nature of their relationship?

Percentage change of Petrol barrels

"\frac{600-500}{500}\times 100 =20" %


Percentage change in price of scooter

"\frac{25000-22000}{25000}\times 100 = 12" %


The cross elasticity of demand

"E_{c} = \frac{20}{12}=1.67"

The goods are substitutes.